Circular No. 8950

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

POSSIBLE NOVA IN OPHIUCHUS S. Nakano, Sumoto, Japan, reports the discovery of another possible

nova (mag 11.3) in Ophiuchus by K. Nishiyama (Kurume, Fukuoka-ken, Japan) and F. Kabashima (Miyaki-cho, Saga-ken, Japan) on two unfiltered 30-s CCD frames (limiting mag 13.6) taken on May 31.608 and 31.609 UT using a 105-mm f/5.6 camera lens. Their confirming unfiltered CCD images (limiting mag 18.5) with a 40-cm reflector taken around May 31.628 yield the following precise position: α = 17^{h}33^{m}29^{s}.67, δ = −27^{o}01′16′′.4 (equinox 2000.0); their image from May 31.795 yields position end figures 29^{s}.64, 15′′.9. They add that nothing is visible at this position on Digitized Sky Survey red images. NSV 22564 is located some 2′ away. Following posting of this object on the Central Bureau’s unconfirmed-objects webpage, several other observers have obtained confirming photometry and astrometry of the new variable. Additional magnitudes for this possible nova, from unfiltered CCD images unless otherwise noted: 1985 May 13, [19.5 (Digitized Sky Survey red plate taken at the European Observatory, via C. Jacques and E. Pimentel, Belo Horizonte, Brazil); 1991 Aug. 11, [20: (U.K. Schmidt Telescope red plate, via E. Guido and G. Sostero, Remanzacco, Italy); 2008 May 11.661, [13.6 (Nishiyama and Kabashima); 20.649, [12.6 (Nishiyama and Kabashima); 25.688, [11.8 (Nishiyama and Kabashima); 31.765, 11.1 (K. Kadota, Ageo, Japan, 25-cm reflector); 31.795, 11.4 (Nishiyama and Kabashima); June 1.540 and 1.595, 11.7 (K. Itagaki, Teppo-cho, Yamagataken, Japan, 60-cm reflector); 1.576, 11.3 (Kadota); 2.022, R ≃ 11.7 (H. Mikuž, Crni Vrh Observatory, 19-cm reflector); 2.23, B = 13.9, V = 12.8, R = 11.4 (Guido and Sostero, remotely with a 25-cm reflector near Mayhill, New Mexico); 2.39, V = 12.8 (Jacques and Pimentel, remotely with a 25-cm reflector near Mayhill). All of the non-discoverer observers above have also provided astrometry for this apparent nova, yielding position end figures 29^{s}.61 ± 0^{s}.01, 14′′.5 ± 0′′.3. Guido and Sostero have posted an image of this transient at the following URL: http://tinyurl.com/4vj2qz.

COMET P/2008 J3 (McNAUGHT) Additional observations have led to improved orbital elements that

show this comet (cf. IAUC 8942) to be of short period: T = 2009 Mar. 11.043 TT, q = 2.27346 AU, e = 0.41559, ω = 5^{o}.035, Ω = 9^{o}.764, i = 25^{o}.400 (equinox 2000.0), P = 7.67 years (from MPEC 2008-K54).

2008 June 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green